Feb. 25-28

– Four stores provided ~33.35 hours of luminosity
– NuMI target cooling leak investigation continues
– Tevatron personnel conducted end-of-store studies
– Linac personnel repairs low-energy buncher
– Main Injector resolved power supply phasing problems
– DZero required eight-hour access for repairs
– Store 8543 lost due to lightning strike
– Shot setup halted due to RF trip and quench

The integrated luminosity for the period from Feb. 21-28 was 41.81 inverse picobarns. NuMI reported receiving 3.98E18 protons on target during this same period.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts